OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the first equation is substituted into the second equation. 12x – y = –4 4x – 3y = –6

OpenStudy (vishweshshrimali5):

4x – 3y = –6 => 12 x - 9 y = -18 now 12x – y = –4 12 x - 9 y = -18 _____________________ - + + 0 + 8y = 14 => y = 7/4 now put this value in 4x - 3y = -6 => 4x - 21/4 = -6 => 4x = -3/4 => x = -3/16